# BenTo: Benchmark Task Reduction with In-Context Transferability

### v1

@@ 1 Introduction
Table 1: Our evaluation of LLMs on two benchmarks and their BenTo-reduced versions. Publicly reported results are available in Appendix G.
@@
To effectively extract a representative subset of tasks that mirrors the full scope of the original benchmark, we propose Benchmark Task reductiOn (BenTo) that formulates the task selection into a facility location (FL) problem. In BenTo, the task similarities are derived either directly from the similarity matrix computed via Laplacian Eigenmaps (LE) or are recalculated within the LE-embedded space. The FL objective was to maximize the similarity between each task in the benchmark and the closest task in the reduced subset. This objective is submodular, allowing us to employ a greedy algorithm that efficiently achieves a high-quality approximate solution. Extensive experiments are conducted to evaluate the effectiveness of BenTo-reduced benchmark by comparing the performance of several widely used LLMs on both the reduced and original benchmarks. Remarkably, as is shown in Footnote 2 (RIGHT) and Table 1, the results are highly consistent, even though the reduced benchmark comprises only 5% of the original tasks. This finding underscores the efficiency of our approach. When compared to existing benchmark reduction methods, BenTo not only yields more accurate evaluation results but also significantly lowers the costs associated with transferability estimation. Furthermore, ICT offers substantial potential benefits across a wide array of LLM research problems and applications, making it a topic of independent interest.
@@ 2 Related Work
Task Transferability. Efficient estimation of task transferability has been a long-studied problem. Past works mainly use Bayesian optimization (Weiss et al., 2016) and information theory (Bao et al., 2019; Tan et al., 2021). LEEP (Nguyen et al., 2020) proposes estimating via approximately training the model by linear probing on the data of the source tasks and evaluate on the target tasks, which resembles our in-context learning approach. Xia et al. (2024) leverages similarity between gradient features obtained by training with LoRA (Hu et al., 2022) as a transferability measure, which inspires us to transform the performance feature into similarity matrices.
Benchmark Reduction. Current benchmark reduction methods can be categorized into two major approaches: selecting tasks from a benchmark and selecting examples from a single task. Our work falls into the first one. It may seem that the second approach is more robust at least on non-few-shot benchmarks (Perlitz et al., 2024), but we can prove that combining the two approaches always yields a better result (see Section 5.2), so these two approaches are actually parallel and we can apply them one by one. Ye et al. (2023) also goes in the first direction and analyzes Big-bench (bench authors, 2023). However, it’s time-consuming to collect the training data they need, as their methods require performances from different models with various parameters as guidance. Polo et al. (2024) goes in the second direction, making use of item response theory (IRT) models, but share the same drawback. On the contrary, our method only requires the performance of a single model as the guidance, making it very efficient in data collection. Vivek et al. (2024) is a relatively efficient example-selection method, which, similar to us, also draws insight from clustering. Yet, their approach can only predict the ranking instead of the exact performance of models on the whole benchmark.
@@ 3 Task Transferability Analysis by In-Context Learning
Aiming at proposing a cost-efficient benchmark reduction approach, we first analyze the structure of benchmarking tasks by studying the transferability between each two tasks. Given source task-$i$ and target task-$j$, the transferbility from task-$i$ to task-$j$ is often measured by training a model on task-$i$ then evaluating it on task-$j$. However, this approach requires training per task and thus is not computationally scalable to modern LLMs and many tasks. We propose to harness in-context learning to provide a training-free estimation of the transferability between tasks.
@@
To study the structure of a multi-task benchmark, we estimate an matrix of task transferability for all the pairs of tasks by applying the above operation to each pair. Assuming that we have $N$ tasks in total, then we will get an $N\times N$ transferability matrix $A$, where $A_{ij}$ is an estimation of the in-context transferability (ICT) from task-$i$ to task-$j$, i.e.,
@@
where $n_{j}$ is the total number of input-output pairs we sampled from task-$j$. $s(\cdot,\cdot)$ is an evaluation metric such as an exact match or similarity score. A larger $s(\cdot,\cdot)$ indicates a better transfer ICL’s performance. To further reduce variance, we resample the $n_{j}$ questions for multiple times using different random seeds and average the achieved $A_{ij}$. Since target tasks may differ in difficulty, we normalize $A$ by zero-centering each column of $A$, i.e.,
@@
Spectral clustering. We investigate the graph structure among tasks by applying clustering based on the tasks’ feature matrix $A$, which can reveal whether the transferability between intra-cluster tasks is high (and thus they can be further reduced). Since $A$ defines the pairwise transferability on a graph of tasks, we choose to use spectral clustering, a widely-used algorithm for graph cut problems. Given $A$, spectral clustering computes a similarity matrix that is symmetric and non-negative, by applying a Euclidean similarity kernel to $A$:
@@
where $c$ is a constant to ensure the non-negativeness of $S$. We discuss the choice of $c$ in Section 5.2. With a similarity matrix $S$ given, spectral clustering can be performed as shown in Footnote 2, where each cluster is defined by a very clear theme. For example, the red cluster contains all three tasks about history in the MMLU benchmark. The clustering result aligns well with human intuitions, which demonstrates the effectiveness of ICT on capturing the inter-task transferability and redundancy between benchmarking tasks. Note that there are a few tasks with counter-intuitive clustering assignments, e.g., “professonal law” is in the biology cluster. This may indicate that our ICT representation capture information inherent to the task structures, which cannot be inferred solely from their names. We will come back to this in Section 5.
@@
The clustering results motivate us to select tasks that are more important or representative: these tasks exhibit higher transferability to more tasks than others, and each of them may transfer to a different subset of tasks. If we can select a subset of representative tasks that can cover most of tasks in a benchmark, then we can use the performance on this subset to approximate the performance on the whole benchmark! That is the benchmark-task reduction problem we’ll discuss in the next section.
@@ 4 Benchmark-Task Reduction (BenTo) by Facility Location
Benchmark-task reduction (BenTo) aims to select a fixed-size subset of tasks, such that the performance of any model on this subset serves as a proxy of the model’s performance on the whole benchmark. Intuitively, we want to choose the most “representative” subsets of tasks. In the context of transferability, the representativity of a subset of tasks can be measured by their transferability to other tasks in the benchmark. If a subset can transfer to the whole benchmark well, then the performance on this subset will be highly correlated to the performance on the whole benchmark. Under the assumption that the task difficulty in a benchmark is on similar levels, the model performance on this subset can directly serve as a prediction of the performance on the whole benchmark, and this holds for any models.
How do we choose such a subset? For every target task in the benchmark, if we can always find a source task in our subset with sufficiently high transferability to the target, then the subset can transfer to all the tasks in the the benchmark. On the other hand, if there exist two or more tasks with high transferability to the same group of tasks, then retaining only one of them suffices to keep the transferability of the representative subset. Inspired by these intuitions, we aim to find a subset such that the similarity between each task in the benchmark and its “nearest” task in the subset is maximized. This formally reduces to optimizing the facility location (FL) function, i.e.,
@@
As shown in Figure 2, $S^{\prime}$ share a lot of common properties with $S$, so we propose a variant that replaces $S$ in Equation 5 with $S^{\prime}$. The LE embedding removes some long-range noise and is expected to be more robust in realistic scenarios. We call the original version using $S$BenTo-sim” and the LE variant using $S^{\prime}$enTo-le”, where BenTo is an abbreviation of Benchmark Task reductiOn.
@@ 5 Experiment
Benchmarks. We assess our method mainly on two benchmarks: MMLU (Hendrycks et al., 2021a; b) and FLAN (Wei et al., 2021). Additional results on AGIEval English (Zhong et al., 2023) and Big-Bench Hard (Suzgun et al., 2022) can be found in Appendix F. MMLU is a question-answering dataset containing 57 tasks with diverse subjects and levels of difficulty, mainly focus on the knowledge of the language models. All the questions in MMLU are multiple-choice questions with 4 options for each answer. On MMLU, we use accuracy (ACC) as the evaluation metric $s(\cdot,\cdot)$ in Equation 1. FLAN is a dataset with more diverse forms of tasks, including free-form generation tasks like translation and summarization. Since FLAN is a large dataset, we sampled 100 questions from each of its 66 tasks as our “whole benchmark”. On FLAN, we use response perplexity as the evaluation metric $s(\cdot,\cdot)$. We follow widely used prompts for these benchmarks without any re-engineering, more details in Appendix D.
@@ 5.2 Ablation Study
Facility location v.s. K-medoids clustering. We select FL for its efficiency and precise formation of the task reduction problem. To evaluate this choice, we compare FL with more computationally intensive methods like K-medoids clustering, which can be viewed as K-means where real data points serve as centroids. As illustrated in Figure 4, the choice of ICL embedding plays a far more critical role in determining performance than the choice of clustering algorithm. When using the same embedding, both FL and K-medoids produce comparable results; however, FL offers a clear advantage in computational efficiency. This makes FL the more practical option for large-scale scenarios without sacrificing performance.
@@ Appendix B Experiment details
For our main experiments, we use 4 A100 40G for about 3 days. For a few models that require a higher RAM, we use 4 H100 for less than 2 hours. We use $L=5$ exemplars and $M=10$ random seeds. We set $Q$ to be a large value so that we always evaluate on the whole test set. For our main experiments on FLAN, when we normalize $E$, we also divide it by the standard deviation since the metric we use makes the original one too distorted.
@@ Appendix E Detailed performance of each model
The detailed performance of each model is shown in Table 8. We can see that our reduced benchmark works decently for every model.
@@ Appendix F Results on additional benchmarks
Table 10: NRMSE on Big Bench Hard (lower the better). $k$ is the number of selected tasks. Each number is averaged over 5 different models.

### v3

@@ 1 Introduction
Table 1: Evaluation of LLMs on two benchmarks and their BenTo-reduced versions using the same prompts and random seeds55footnotemark: 5. Previously reported results [^6] The evaluation error on BBH is higher due to its higher variance in ICL evaluations. This is also reflected by the larger difference between our evaluation and previously reported results in Appendix G.are available in Appendix G.
@@
66footnotetext: They may use slightly different prompts and distinct random seeds, which are not released.
To effectively extract a representative subset of tasks that mirrors the full scope of the original benchmark, we propose Benchmark Task reductiOn (BenTo) that formulates the task selection into a facility location (FL) problem (Cornuejols et al., 1977). In BenTo, the task similarities are derived either directly from the similarity matrix computed via Laplacian Eigenmaps (LE) or are recalculated within the LE-embedded space. The FL objective was to maximize the similarity between each task in the benchmark and the closest task in the reduced subset. This objective is submodular, allowing us to employ a greedy algorithm (Nemhauser et al., 1978) that efficiently achieves a high-quality approximate solution. Extensive experiments are conducted to evaluate the effectiveness of BenTo-reduced benchmark by comparing the performance of several widely used LLMs on both the reduced and original benchmarks. Remarkably, as is shown in Footnote 2 (RIGHT) and Table 1, the results are highly consistent, even though the reduced benchmark comprises only 5% of the original tasks. This finding underscores the efficiency of our approach. When compared to existing benchmark reduction methods, BenTo not only yields more accurate evaluation results but also significantly lowers the costs associated with transferability estimation. Furthermore, ICT offers substantial potential benefits across a wide array of LLM research problems and applications, making it a topic of independent interest.
@@ 2 Related Work
Task Transferability. Efficient estimation of task transferability has been a long-studied problem. Past works mainly use Bayesian optimization (Weiss et al., 2016) and information theory (Bao et al., 2019; Tan et al., 2021). LEEP (Nguyen et al., 2020) proposes estimating via approximately training the model by linear probing on the data of the source tasks and evaluating the target tasks, which resembles our in-context learning approach. Xia et al. (2024) leverages similarity between gradient features obtained by training with LoRA (Hu et al., 2022) as a transferability measure, which inspires us to transform the performance feature into similarity matrices.
Benchmark Reduction. Dataset reduction for LLM training Li et al. (2024b; a) has been a heated area while the reduction for benchmarks is till under-explored. Current benchmark reduction methods can be categorized into two major approaches: selecting tasks from a benchmark and selecting examples from a single task. Our work falls into the first one. It may seem that the second approach is more robust at least on non-few-shot benchmarks (Perlitz et al., 2024), but we can prove that combining the two approaches always yields a better result (see Section 5.2), so these two approaches parallel and we can apply them one by one. Ye et al. (2023) also goes in the first direction and analyzes Big-bench (bench authors, 2023). However, it’s time-consuming to collect the training data they need, as their methods require performances from different models with various parameters as guidance. Polo et al. (2024) goes in the second direction, making use of item response theory (IRT) models, but share the same drawback. On the contrary, our method only requires the performance of a single model as guidance, making it very efficient in data collection. Vivek et al. (2024) is a relatively efficient example-selection method, which, similar to us, also draws insight from clustering. Yet, their approach can only predict the ranking instead of the exact performance of models on the whole benchmark.
@@ 3 Task Transferability Analysis by In-Context Learning
Aiming at proposing a cost-efficient benchmark reduction approach, we first analyze the structure of benchmarking tasks by studying the transferability between each pair of tasks. Given source task-$i$ and target task-$j$, the transferability from task-$i$ to task-$j$ is often measured by training a model on task-$i$ and then evaluating it on task-$j$. However, this approach requires training per task and thus is not computationally scalable to modern LLMs and many tasks. We propose to harness in-context learning to provide a training-free estimation of the transferability between tasks.
@@
To study the structure of a multi-task benchmark, we estimate a matrix of task transferability for all the pairs of tasks by applying the above operation to each pair. Assuming that we have $N$ tasks in total, then we will get an $N\times N$ transferability matrix $A$, where $A_{ij}$ is an estimation of the in-context transferability (ICT) from task-$i$ to task-$j$, i.e.,
@@
where $n_{j}$ is the total number of input-output pairs we sampled from task-$j$. $s(\cdot,\cdot)$ is an evaluation metric such as an exact match or similarity score. A larger $s(\cdot,\cdot)$ indicates a better transfer-ICL’s performance. To further reduce variance, we resample the $n_{j}$ questions multiple times using different random seeds and average the achieved $A_{ij}$. Since target tasks may differ in difficulty, we normalize $A$ by zero-centering each column of $A$, i.e.,
@@
Spectral clustering. We investigate the graph structure among tasks by applying clustering based on the tasks’ feature matrix $A$, which can reveal whether the transferability between intra-cluster tasks is high (and thus they can be further reduced). Since $A$ defines the pairwise transferability on a graph of tasks, we choose to use spectral clustering, a widely used algorithm for graph cut problems. Given $A$, spectral clustering computes a similarity matrix that is symmetric and non-negative, by applying a Euclidean similarity kernel to $A$:
@@
where $c$ is a constant to ensure the non-negativeness of $S$. We discuss the choice of $c$ in Section 5.2. With a similarity matrix $S$ given, spectral clustering can be performed as shown in Footnote 2, where each cluster is defined by a very clear theme. For example, the red cluster contains all three tasks about history in the MMLU benchmark. The clustering result aligns well with human intuitions, which demonstrates the effectiveness of ICT on capturing the inter-task transferability and redundancy between benchmarking tasks. Note that there are a few tasks with counter-intuitive clustering assignments, e.g., “professional law” is in the biology cluster. This may indicate that our ICT representation captures information inherent to the task structures, which cannot be inferred solely from their names. We will come back to this in Section 5.
@@
The clustering results motivate us to select tasks that are more important or representative: these tasks exhibit higher transferability to more tasks than others, and each of them may transfer to a different subset of tasks. If we can select a subset of representative tasks that can cover most of the tasks in a benchmark, then we can use the performance on this subset to approximate the performance on the whole benchmark! That is the benchmark-task reduction problem we’ll discuss in the next section.
@@ 4 Benchmark-Task Reduction (BenTo) by Facility Location
Benchmark-task reduction (BenTo) aims to select a fixed-size subset of tasks, such that the performance of any model on this subset serves as a proxy of the model’s performance on the whole benchmark. Intuitively, we want to choose the most “representative” subsets of tasks. In the context of transferability, the representativity of a subset of tasks can be measured by their transferability to other tasks in the benchmark. If a subset can transfer to the whole benchmark well, then the performance on this subset will be highly correlated to the performance on the whole benchmark. Under the assumption that the task difficulty in a benchmark is on similar levels, the model performance on this subset can directly serve as a prediction of the performance on the whole benchmark, and this holds for any model.
How do we choose such a subset? For every target task in the benchmark, if we can always find a source task in our subset with sufficiently high transferability to the target, then the subset can transfer to all the tasks in the benchmark. On the other hand, if there exist two or more tasks with high transferability to the same group of tasks, then retaining only one of them suffices to keep the transferability of the representative subset. Inspired by these intuitions, we aim to find a subset such that the similarity between each task in the benchmark and its “nearest” task in the subset is maximized. This formally reduces to optimizing the facility location (FL) function, i.e.,
@@
As shown in Figure 2, $S^{\prime}$ shares a lot of common properties with $S$, so we propose a variant that replaces $S$ in Equation 5 with $S^{\prime}$. The LE embedding removes some long-range noise and is expected to be more robust in realistic scenarios. We call the original version using $S$BenTo-sim” and the LE variant using $S^{\prime}$enTo-le”, where BenTo is an abbreviation of Benchmark Task reductiOn.
@@ 5 Experiment
Benchmarks. We assess our method mainly on two benchmarks: MMLU (Hendrycks et al., 2021a; b) and FLAN (Wei et al., 2021). Additional results on AGIEval English (Zhong et al., 2023) and Big-Bench Hard (Suzgun et al., 2022) can be found in Appendix F. MMLU is a question-answering dataset containing 57 tasks with diverse subjects and levels of difficulty, mainly focusing on the knowledge of the language models. All the questions in MMLU are multiple-choice questions with 4 options for each answer. On MMLU, we use accuracy (ACC) as the evaluation metric $s(\cdot,\cdot)$ in Equation 1. FLAN is a dataset with more diverse forms of tasks, including free-form generation tasks like translation and summarization. Since FLAN is a large dataset, we sampled 100 questions from each of its 66 tasks as our “whole benchmark”. On FLAN, we use response perplexity as the evaluation metric $s(\cdot,\cdot)$. We follow widely used prompts for these benchmarks without any re-engineering, more details in Appendix D.
@@ 5.2 Ablation Study
Facility location v.s. K-medoids clustering. We select FL for its efficiency and precise formation of the task reduction problem. To evaluate this choice, we compare FL with more computationally intensive methods like K-medoid clustering, which can be viewed as K-means where real data points serve as centroids. As illustrated in Figure 4, the choice of ICL embedding plays a far more critical role in determining performance than the choice of clustering algorithm. When using the same embedding, both FL and K-medoids produce comparable results; however, FL offers a clear advantage in computational efficiency. This makes FL the more practical option for large-scale scenarios without sacrificing performance.
@@ Appendix B Experiment details
For our main experiments, we use 4 A100 40G for about 3 days. We use $L=5$ exemplars and $M=10$ random seeds. We set $Q$ to be a large value so that we always evaluate on the whole test set. For our main experiments on FLAN, when we normalize $E$, we also divide it by the standard deviation since the metric we use makes the original one too distorted.
@@ Appendix E Detailed performance of each model
The detailed performance of each model is shown in Table 8. Our reduced benchmark works consistently well across different models.
@@ Appendix F Results on additional benchmarks
Table 10: NRMSE on Big Bench Hard (lower the better). $k$ is the number of selected tasks. Each number is averaged over 8 different models.